# Pay Less Attention to Function Words for Free Robustness of Vision-Language Models

### v2

@@ Abstract
To address the trade-off between robustness and performance for robust VLM, we observe that function words could incur vulnerability of VLMs against cross-modal adversarial attacks, and propose Function-word De-Attention (FDA) accordingly to mitigate the impact of function words. Similar to differential amplifiers, our FDA calculates the original and the function-word cross-attention within attention heads, and differentially subtracts the latter from the former for more aligned and robust VLMs. Comprehensive experiments include 2 SOTA baselines under 6 different attacks on 2 downstream tasks, 3 datasets, and 3 models. Overall, our FDA yields an average 18/13/53% ASR drop with only 0.2/0.3/0.6% performance drops on the 3 tested models on retrieval, and a 90% ASR drop with a 0.3% performance gain on visual grounding. We demonstrate the scalability, generalization, and zero-shot performance of FDA experimentally, as well as in-depth ablation studies and analysis. Code will be made publicly available.
@@ 1 Introduction
Building robust vision-language models (VLMs) has gathered profound academic focus because of the necessity of defending VLMs against various adversarial attacks. To this end, many works [17, 13] have been proposed to enhance model robustness, purify perturbations, or detect potential adversaries. Among them, adversarial training (AT) shows superior performance in enhancing the robustness of VLMs. However, AT-based methods incur significant performance drops compared to vanilla models and high computational costs.
To resolve the trade-off mentioned above, we propose to enhance VLM robustness by further refining vision-language alignment (VLA). Rather than perturbing images during fine-tuning, we break texts into finer grains: content words, i.e., nouns/verbs, and function words, i.e., am/is/are. Specifically, we hypothesize that function words could incur vulnerability of VLMs against cross-modal adversarial attacks because of their ubiquity and lack of specificity. To verify our hypothesis, we record the white-box similarity between function-/content-words and images during targeted (image) attacks [^1] We tested on the 1k testset of Flickr30k retrieval dataset, using PGD attack with $\epsilon={4}/{255}$., and find that 80.3% of images show higher similarity scores towards the function words than content words after attacks, while 0% of the images exhibit this pattern before attacks. We also provide a visual demonstration using Grad-CAM[18] from a successful untargeted attack to demonstrate the impact of function words. As shown in Fig.1, removing all function words greatly mitigates the distractions from adversarial perturbations. Lastly, to qualitatively validate the impact of function words in adversarial examples, we record the performance variation after removing nouns, adjectives, verbs, and function words on both clean and adversarial examples for Image-to-Text/Text-to-Image Retrieval (T2IR/I2TR), as presented in Table.1. Results show that function words are the only words that reduce ASR without causing a significant performance drop. These results confirm our hypothesis on function words, implying that proper removal of function words could potentially defend VLMs against attacks.
@@
Consequently, inspired by the setting of differential transformers [22] and differential amplifiers, we propose Function-word De-Attention (FDA) as the first method to build robust VLMs by refining vision-language alignment. Specifically, our FDA works by deploying a parallel pipeline on multi-attention heads within fusion-encoders, calculating the cross-attention between function words and the input images, i.e., distractions. We further softmax along the dimensions of visual and textual tokens to highlight the most misleading textual/visual tokens. Finally, we subtract the above distractions from the original attention for the output. To validate the effectiveness of FDA, we conduct comprehensive experiments on two SOTA baselines, 3 models, 2 tasks, 3 datasets, and 6 attacks. Overall, our FDA yields an average 18/13/53% ASR drop with only 0.2/0.3/0.6% performance drops on the 3 tested models for retrieval, and a 90% ASR drop with better clean performance on grounding. Our FDA is also verified to enhance the zero-shot performance of VLMs.
Table 1: Effectiveness of removing different words when testing on clean and adversarial examples. Adversarial examples use AutoAttack, and $\Delta_{ASR}$ is presented using the average results for all epsilons ($\epsilon=2,4$). $\uparrow$ $\downarrow$ indicates increased/decreased $\Delta_{ASR}$ (higher values preferred). Removing function words can lower ASR without significantly harming clean performance.
Removed Clean (R@1) ($\uparrow$) Avg ASR Drop
Words T2IR I2TR $\Delta_{ASR}$ ($\uparrow$)
N/A 95.90 85.60 -
NOUN 58.90 32.83 $\uparrow$ 25.27
ADJ 91.60 78.36 $\downarrow$   0.42
VERB 93.80 79.36 $\downarrow$   0.38
FUNC 94.30 81.04 $\uparrow$   0.54
Figure 1: Grad-CAM of attention maps of VLM under white-box untargeted attacks through perturbed images. The texts are given at the bottom of the figure, with function words highlighted. Left: The VLM correctly recognizes the female student on the clean image given the token her. Mid: The VLM is distracted by the adversarial perturbation and partially looks at the male coach. Right: The distraction is mitigated by simply applying masks to remove all function words: the VLM successfully ‘looks back at’ the female student.
@@
  • We identify that function words are distractions for vision-language alignment and subsequently propose Function-word De-Attentioning (FDA) to pay less attention to function words for more aligned vision-language models with free robustness.
  • We conduct comprehensive experiments on two SOTA baselines, 3 models, 2 tasks, and 3 datasets, under 6 attacks, and validate the effectiveness of FDA in enhancing robustness while preserving performance.
  • We provide in-depth ablation studies to show qualitative analysis, the insensitivity of our FDA towards hyperparameters, generalization across backbones, and enhancement on zero-shot performance.
@@ 2 Related Work
Adversarial attacks on vision-language models. In light of the advancement in VLMs, adversarial attacks on VLMs have also emerged to fool VLMs into incorrect or misleading outputs. Recent studies on white-box attacks [1] have exhibited impressive results. Besides, black-box attacks[26, 11, 23, 5, 20] have also demonstrated significant effectiveness against pre-trained VLMs through transferable attacks.
Adversarial Defense on vision-language models For defenses, adversarial training (AT) [16, 25, 19] has exhibited significant effectiveness in defending models against various adversarial attacks against classification, retrieval, etc. Several AT-based methods [17, 13] have demonstrated impressive robustness boost on CLIP models. However, AT is notoriously well-known for downgrading performance significantly due to the inclusion of adversarial examples into training. Although [17] proposed FARE to use the visual embeddings of vanilla models as supervision to balance the trade-off between clean performance and robustness, the performance drops remain considerably noticeable. Besides, the high computational costs also hinder broader applications in practice.
@@ 3.1 Preliminary
For a given textual/visual encoder $\mathcal{T}/\mathcal{V}$, input images $I$ and texts $T$ are fed into repsective encoders with their attention masks $\mathcal{M}_{T}/\mathcal{M}_{I}$ for the embeddings $\mathbf{\mathcal{F}}_{T},\mathbf{\mathcal{F}}_{V}\in\mathbb{R}^{d_{k}}$:
@@ 3.2 Function-word De-Attention (FDA)
Figure 2: Left: An illustration of our Function-word De-Attention (FDA) method. On the existing process of attention calculation, which uses $\mathcal{F}_{V}$ and $\mathcal{F}_{T}$, we add a parallel pipeline to calculate the attentions between function words $\mathcal{F}_{T_{f}}$ and the images $\mathcal{F}_{V}$. Afterwards, the function-attention passes a control gate $\mathcal{G}$ before entering the FDA module (triangle) differentially to subtract distractions as presented in Eq.6. Right: We speculate that attacks can easily cross the boundary for misalignments for less aligned models (top), and by removing function-word distractions, models can learn a robust embedding (bottom), preventing misalignments.
Built upon our previous observation, we hypothesize that function words are potential distractions in vision-language alignment. To remove such distractions, we propose Function-word De-Attention (FDA): we add a parallel pipeline upon the existing cross-attention calculation to specifically acquire the cross-attention between function words and the input images, namely the distraction, and then subtract them from the original attention. An illustration of our FDA is given in Fig.2. We first parallelly extract the features of all function words (denoted as $T_{f}$) within the input texts by masking all other tokens, excluding [CLS]:
@@
where $M_{t_{f}}$ is the function word mask, and $\mathcal{D}$ is the function word dictionary. Here, we use a dictionary shortlisted from the stopwords list in [9]. Subsequently, we adopt a parallel pipeline to calculate function words’ attention scores:
@@
$\tilde{A}tt_{t}^{L,H}=softmax\big(S_{T_{f}}^{L,H},dim=-1\big)V,$ (5)
$\tilde{A}tt_{v}^{L,H}=softmax\big(S_{T_{f}}^{L,H},dim=-2\big)V$
Afterwards, we subtract both distractions from $Att$ individually and take the minimum value as the final attention scores, with $\mathcal{G}$ being a control gate for automatically adjusting the subtraction.
$\hat{A}tt^{L,H}=min\big(Att^{L,H}-\mathcal{G}(\tilde{A}tt_{t}^{L,H}),Att^{L,H}-\mathcal{G}(\tilde{A}tt_{v}^{L,H})\big)$ (6)
Finally, after complete calculation of FDA, denoted as $FDA(\cdot)$, we concatenate the attention scores from all attention heads for outputs:
$\mathbf{\hat{A}tt}^{L,H}=$ (7)
@@
FDA can be flexibly implemented on any number of fusion layers/attention heads, as each may specialize differently [6]. A general intuition is to remove these distractions in the early layers instead of the later ones to avoid possible ‘absorption’ of distractions, but not so exquisitely upfront that it may undermine the contextual integrity of the original inputs. We provide ablation and analysis in Sec.4.3.
@@ 4 Experiments
Vision-Language Tasks&Datasets. To thoroughly evaluate the performance and robustness of FDA, we incorporate several downstream tasks, including Text-to-Image/Image-to-Text Retrieval (T2IR/I2TR) and Visual Grounding (VG). For datasets, we use the Flickr30k[14] and MSCOCO [10] dataset for retrieval, and RefCOCO+ [24] for VG.
@@
Models. For T2IR/I2TR, we test our method on the ALBEF [7], TCL [21], and BLIP[8], using 14M/14M/124M pretrained images, respectively. All models use the ViT-B/32 [3] as visual encoders and BERT [2] as textual encoders. Specifically, TCL shares the same backbones as ALBEF but uses a different training strategy (triplet contrastive learning), while BLIP uses a larger pre-trained encoder with an extra decoder. CLIP [15] is not included due to the absence of fusion encoders.
@@
Baselines. As for baselines, we adopt the two SOTA methods for robust CLIP, i.e., TeCoA [13] and FARE [17], on all the models as adversarial fine-tuning baselines. To account for the robustness and accuracy trade-off, we lower the perturbation strength of each method to ensure similar clean performance as our FDA, such that TeCoA and FARE serve as a reference to compare robustness. Specifically, we use $\epsilon=1$ for TeCoA and FARE. For Text-to-Image/Image-to-Text Retrieval, we adversarially fine-tune using TeCoA and FARE for 4/1/1 epochs for ALBEF/TCL/BLIP. For Visual Grounding, we adversarially train models with both TeCoA and FARE for only 1 epoch, as both methods incur significant performance drops on ALBEFs.
Attacks. To thoroughly evaluate the robustness of our models, we test all models with three adversarial attacks and use the average of all attacks for robustness evaluations. Specifically, we use Projected Gradient Descent [12] and AutoAttack [1], denoted as PGD and APGD. As for the adaptive attack, we apply function word masks to the input texts for APGD to evade our FDA, denoted Masked APGD (MAPGD). All attacks are fully white-box, i.e., the attackers are aware of and can access the extra FDA operations. For each attack, we follow the settings in [13] and [17] to attack images with perturbation bounded by $l_{\infty}=\frac{2}{255},\frac{4}{255}$. Specifically, for targeted attacks in T2IR/I2TR, we apply a circular shift targeted to ensure non-overlapping unmatched targets. For targeted attacks in VG, we follow the settings of [4] and do not apply a patched attack.
Metrics. We use the common metric, Attack Success Rate (ASR), to indicate the efficacy of all adversarial attacks. For an ASR on a given model and the baseline model, denoted as $ASR_{M}/ASR_{B}$, respectively, we calculate the relative ASR change in percentage using $\Delta_{ASR}=\frac{ASR_{B}-ASR_{M}}{ASR_{B}}\times 100\%$. Consequently, a positive/larger $\Delta_{ASR}$ indicates improved/stronger robustness, while a negative/lower $\Delta_{ASR}$ implies decreased/weaker robustness, with 0% (100%) meaning no robustness gain (completely defended). Details are given in the Sec.A of the Appendix.
Implementation Details. Since our FDA parallelly computes distracted attention for subtraction, finetuning models with FDA is identical to downstream finetuning without extra modifications or parameters. Following the settings in [7], we finetune the model by 10 epochs and use the last-epoch model for all tasks/models. For the layer index, we use $L^{0-1}$ and $H^{0-5}$ for all models/tasks/datasets, with corresponding ablation studies on the selection of the layer and attention head indices for all models and tasks in Sec.4.3.
@@
Table 2: Attack success rate (ASR) of PGD/APGD/MAPGD (masked APGD) against for Text-to-Image/Image-to-Text Retrieval (T2IR/I2TR) on Flickr30k and $CO$. Results are presented in percentage (%). $\uparrow$ $\downarrow$ indicates increased/decreased $\Delta_{ASR}$ (higher values preferred). $\dagger$ indicates higher performance than clean models. (Full results are given in Sec.B of the Appendix.) Our FDA consistently shows the best performance and overall robustness on ALL models.
@@
BLIP
$\nicefrac{{2}}{{255}}$ No Defense 97.20 25.10 63.26 50.19 87.30 11.83 60.08 44.35 -
@@
Table 3: Attack success rate (ASR) of PGD/APGD/MAPGD (masked APGD) against for Visual Grounding (VG) on RefCOCO+. Results are presented in percentage (%). $\uparrow$ $\downarrow$ indicates increased/decreased $\Delta_{ASR}$ (higher values preferred). $\dagger$ indicates higher performance than clean models. (Full results are given in Sec.B of the Appendix.) Our FDA consistently shows the best performance and overall robustness on ALL models.

### v3

@@ Abstract
To address the trade-off between robustness and performance for robust VLM, we observe that function words could incur vulnerability of VLMs against cross-modal adversarial attacks, and propose Function-word De-Attention (FDA) accordingly to mitigate the vulnerability brought by function words. Inspired by differential transformers, our FDA calculates the original and the function-word cross-attention within attention heads, and differentially subtracts the latter from the former for more robust alignment. Comprehensive experiments include 2 SOTA baselines under 6 different attacks on 2 downstream tasks, 3 datasets, and 3 models. Overall, our FDA yields an average 18/13/53% ASR drop with only 0.2/0.3/0.6% performance drops on the 3 tested models on retrieval, and a 90% ASR drop with a 0.3% performance gain on visual grounding. We demonstrate the scalability, generalization, and zero-shot performance of FDA experimentally, as well as in-depth ablation studies and analysis. Code is available at https://github.com/michaeltian108/FDA.
@@ 1 Introduction
Building robust vision-language models (VLMs) has garnered profound academic focus because of the necessity of defending VLMs against various adversarial attacks. To this end, many works have been proposed to defend models against adversarial images by enhancing model robustness (Schlarmann et al., 2024; Mao et al., 2022), purifying perturbations (Yoon et al., 2021; Lei et al., 2025), or detecting potential adversaries (Metzen et al., 2017; Klingner et al., 2022). Among them, adversarial training (AT) shows superior enhancement in the robustness of VLMs, at cost of significant performance drops and computational costs.
In this paper, we propose to enhance adversarial robustness from an overlooked yet intuitive perspective. We hypothesize that, due to the innate semantic gaps between the clean image and targeted texts, cross-modal perturbation are likely guided towards closer tokens (i.e., smaller semantic gaps). In this regard, function words, which carry little semantic information but have high word frequency, become more approachable than other tokens, incurring unwanted vulnerability. To demonstrate the above hypothesis, we record and visualize the relative clean R@1 and Attack Success Rate (ASR) when removing verbs, adjectives, and function words, compared to not removing. As shown in the left of Fig.1, only function words reduce ASR among all word types and cause the least clean performance drop. The results qualitatively confirms the impact of function words regarding in bringing adversarial vulnerability. We further provide the attention maps (Selvaraju et al., 2017) to visualize the impact of removing different words, presented in the right of Fig.1. All above results validate our above hypothesis, implying that function words carry certain adversarial semantic, and removing them properly could tighten the cross-modal alignment, defending VLMs against attacks.
Figure 1: Left: Histograms on of relative clean R@1 and ASR of No remove, Removing Adj, Verbs and Function words. Results are collected from Flickr30k retrieval testset on ALBEF. Results qualitatively show that removing function words helps reduce ASR and maintain clean performance. Right: Attention maps of VLM under white-box attacks through perturbed images. The texts are given at the bottom, with words marked by corresponding colors. VLM correctly recognizes the corresponding character on the clean image given the token her, but gets distracted towards the incorrect character on the adversarial image. Only removing function words helps the VLM ‘looks back at’ the correct character.
@@
Consequently, inspired by the setting of differential transformers (Ye et al., 2024) and differential amplifiers, we propose Function-word De-Attention (FDA) as the first method to build robust VLMs by refining vision-language alignment. Specifically, our FDA works by deploying a parallel pipeline on multi-attention heads within fusion-encoders, calculating the cross-attention between function words and the input images, i.e., distractions. We further softmax along the dimensions of visual and textual tokens to highlight the most misleading textual/visual tokens. Finally, we subtract the above distractions from the original attention for the output. To validate the effectiveness of FDA, we conduct comprehensive experiments on two SOTA baselines, 3 models, 2 tasks, 3 datasets, and 6 attacks. Overall, our FDA yields an average 18/13/53% ASR drop with only 0.2/0.3/0.6% performance drops on the 3 tested models for retrieval, and a 90% ASR drop with better clean performance on grounding. We also find that FDA improves the zero-shot performance.
@@
1) We identify that function words are vulnerability for VLMs and subsequently propose Function-word De-Attention (FDA) to pay less attention to function words for more aligned vision-language models with free robustness.
2) We conduct comprehensive experiments on two SOTA baselines, 3 models, 2 tasks, and 3 datasets, under 6 attacks, and validate the effectiveness of FDA in enhancing robustness while preserving performance.
3) We provide ablation studies to show the hyperparameters insensitivity, backbone generalization, and zero-shot performance enhancement of our method with in-depth analysis.
@@ 2 Related Work
Adversarial attacks on vision-language models. In light of the advancement in VLMs (Li et al., 2021; 2022; Radford et al., 2021; Bai et al., 2023; Team and others, 2024; Bai et al., 2025; Yang et al., 2025; Liu et al., 2023),several aspects (Li et al., 2025b; a; Zha et al., 2025), such as token compression and broader capabilities, have garnered enormous research interests. Among them, adversarial attacks on VLMs become the most threatening direction as they fool VLMs into incorrect or misleading outputs. Recent studies on white-box attacks (Croce and Hein, 2020) have exhibited impressive results. Besides, black-box attacksZhang et al. (2022); Lu et al. (2023); Yin et al. (2023); He et al. (2023); Tian et al. (2026) have also demonstrated significant effectiveness against pre-trained VLMs through transferable cross-modal attacks.
Adversarial Defense on vision-language models For defenses, adversarial training (AT) (Rice et al., 2020; Zhang et al., 2019; Tian et al., 2024) has exhibited significant effectiveness in defending models against various adversarial attacks against classification, retrieval, etc. Several AT-based methods (Schlarmann et al., 2024; Mao et al., 2022) have demonstrated impressive robustness boost on CLIP models. However, AT is notoriously well-known for downgrading performance significantly due to the inclusion of adversarial examples into training. Although Schlarmann et al. (2024) proposed FARE to use the visual embeddings of vanilla models as supervision to balance the trade-off between clean performance and robustness, the performance drops remain considerably noticeable. Besides, the high computational costs also hinder broader applications in practice.
@@ 3.1 Preliminary
For a given textual/visual encoder $\mathcal{T}/\mathcal{V}$, input images $I$ and texts $T$ are fed into respective encoders with their attention masks $\mathcal{M}_{T}/\mathcal{M}_{I}$ for the embeddings $\mathbf{\mathcal{F}}_{T},\mathbf{\mathcal{F}}_{V}\in\mathbb{R}^{d_{k}}$:
@@ 3.2 Function-word De-Attention (FDA)
Figure 2: Left: An illustration of our Function-word De-Attention (FDA) method. On the existing process of attention calculation, which uses $\mathcal{F}_{V}$ and $\mathcal{F}_{T}$, we add a parallel pipeline to calculate the attentions between function words $\mathcal{F}_{T_{f}}$ and the images $\mathcal{F}_{V}$. Afterwards, the function-attention passes a control gate $\mathcal{G}$ before entering the FDA module (triangle) differentially to subtract distractions as presented in Eq.6. Right: An illustration on how FDA defends attacks. We speculate that attacks can cross the boundary easily to cause misalignments in vanilla models (top), and by removing function-word distractions, models can learn a tightened embedding (bottom), preventing such misalignments.
Built upon our previous hypothesis that function words are potential distractions in vision-language alignment, we propose Function-word De-Attention (FDA) to remove such distractions: we add a parallel pipeline upon the existing cross-attention calculation to specifically acquire the cross-attention between function words and the input images, namely the distraction, and then subtract them from the original attention. An illustration of our FDA is given in Fig.2. We first parallelly extract the features of all function words (denoted as $T_{f}$) within the input texts by masking all other tokens, excluding [CLS]:
@@
where $M_{t_{f}}$ is the function word mask, and $\mathcal{D}$ is the function word dictionary, which is a dictionary shortlisted from the stopwords list in (Li et al., 2020). We then use a parallel pipeline to calculate function words’ attention scores:
@@
$\tilde{A}tt_{t}^{L,H}=softmax\big(S_{T_{f}}^{L,H},dim=-1\big)V,\qquad\quad\\ \tilde{A}tt_{v}^{L,H}=softmax\big(S_{T_{f}}^{L,H},dim=-2\big)V\\$ (5)
Afterwards, we subtract both attention from $Att$ individually and take the minimum value as the final attention scores, with $\mathcal{G}$ being a control gate for subtraction intensity.
$\hat{A}tt^{L,H}=\min\big(Att^{L,H}-\mathcal{G}(\tilde{A}tt_{t}^{L,H}),Att^{L,H}-\mathcal{G}(\tilde{A}tt_{v}^{L,H})\big)$ (6)
Finally, after complete calculation of FDA, denoted as $\mathbb{D}(\cdot)$, we concatenate the attention scores from all attention heads:
@@
FDA can be flexibly implemented on any number of fusion layers/attention heads, as each may specialize differently (Kang et al., 2025). A general intuition is to remove these distractions in the early layers instead of the later ones to avoid possible ‘absorption’, but not so exquisitely upfront that it may undermine the contextual integrity of the textual inputs. We provide ablation and analysis in Sec.4.3.
@@ 4 Experiments
Vision-Language Tasks&Datasets. To thoroughly evaluate the performance and robustness of FDA, we incorporate several downstream tasks, including Text-to-Image/Image-to-Text Retrieval (T2IR/I2TR) and Visual Grounding (VG). For datasets, we use the Flickr30k(Plummer et al., 2015) and MSCOCO (Lin et al., 2014) dataset for retrieval, and RefCOCO+ (Yu et al., 2016) for VG.
@@
Models. For T2IR/I2TR, we test our method on the ALBEF (Li et al., 2021), TCL (Yang et al., 2022), and BLIP(Li et al., 2022), using 14M/14M/124M pretrained images, respectively.For VG, we use ALBEF . All models use the ViT-B/32 (Dosovitskiy et al., 2021) as visual encoders and BERT (Devlin et al., 2019) as textual encoders. Specifically, TCL shares the same backbones as ALBEF but uses a different training strategy (triplet contrastive learning), while BLIP uses a larger pre-trained encoder with an extra decoder. CLIP (Radford et al., 2021) is not included due to the absence of fusion encoders.
@@
Baselines. As for baselines, we adopt the two SOTA methods for robust CLIP, i.e., TeCoA (Mao et al., 2022) and FARE (Schlarmann et al., 2024), on all the models as adversarial fine-tuning baselines. To account for the robustness and accuracy trade-off, we lower the perturbation strength of each method to ensure similar clean performance as our FDA, such that TeCoA and FARE serve as a reference to compare robustness. Specifically, we use $\epsilon=1$ for TeCoA and FARE. For Text-to-Image/Image-to-Text Retrieval, we adversarially fine-tune using TeCoA and FARE for 4/1/1 epochs for ALBEF/TCL/BLIP. For Visual Grounding, we adversarially train models with both TeCoA and FARE for 1 epoch, as both methods incur significant performance drops.
@@
Table 1: Attack success rate (ASR) of PGD/APGD/MAPGD (masked APGD) against for Text-to-Image/Image-to-Text Retrieval (T2IR/I2TR) on Flickr30k and $CO$. Results are presented in percentage (%). $\uparrow$ $\downarrow$ indicates increased/decreased $\Delta_{ASR}$ (higher values preferred). $\dagger$ indicates higher performance than clean models. (Full results are given in Sec.C of the Appendix.) Our FDA consistently shows the best performance and overall robustness on ALL models.
@@
BLIP $\nicefrac{{2}}{{255}}$ No Defense 97.20 25.10 63.26 50.19 87.30 11.83 60.08 44.35 -
@@
Attacks. To thoroughly evaluate the robustness of our models, we test all models with three adversarial attacks and use the average of all attacks for robustness evaluations. Specifically, we use Projected Gradient Descent (Madry et al., 2017) and AutoAttack (Croce and Hein, 2020), denoted as PGD and APGD. As for the adaptive attack, we apply function word masks to the input texts for APGD to evade our FDA, denoted Masked APGD (MAPGD). All attacks are fully white-box, i.e., the attackers can access FDA. For each attack, we follow the settings in Mao et al. (2022) and Schlarmann et al. (2024) to attack images with perturbation bounded by $l_{\infty}=\frac{2}{255},\frac{4}{255}$. Specifically, for targeted attacks in T2IR/I2TR, we apply a circular shift targeted to ensure non-overlapping unmatched targets. For targeted attacks in VG, we follow the settings of Gao et al. (2024) to not apply patched attacks.
Metrics. We use the common metric, Attack Success Rate (ASR), to indicate the efficacy of all adversarial attacks. For an ASR on a defended model and the baseline model, denoted as $ASR_{M}$ and $ASR_{B}$, respectively, we calculate the relative ASR change in percentage using $\Delta_{ASR}=\frac{ASR_{B}-ASR_{M}}{ASR_{B}}\times 100\%$. Consequently, a positive/larger $\Delta_{ASR}$ indicates improved/stronger robustness, while a negative/lower $\Delta_{ASR}$ implies decreased/weaker robustness, with 0% (100%) meaning no robustness gain (completely defended). Details are given in the Sec.B of the Appendix.
Implementation Details. Since our FDA parallelly computes distracted attention for subtraction, fine-tuning models with FDA is identical to downstream fine-tuning without extra modifications or parameters. Following the settings in Li et al. (2021), we fine-tune the model by 10 epochs and use the last-epoch model for all tasks/models. For the layer index, we use $L^{0-1}$ and $H^{0-5}$ for all models/tasks/datasets, with ablation studies on the selection of the layer and attention head indices for all models and tasks in Sec.4.3.
Table 2: Attack success rate (ASR) of PGD/APGD/MAPGD against for Visual Grounding (VG) on RefCOCO+. Results are presented in percentage (%). $\uparrow$ $\downarrow$ indicates increased/decreased $\Delta_{ASR}$ (higher values preferred). $\dagger$ indicates higher performance than clean models. (Full results are given in Sec.C of the Appendix.) Our FDA consistently leads in performance and overall robustness on ALL models.